What price maximizes total expenditures for consumers having a market demand curve P=40-2Q
Answer #1

ANswer

The expenditure is maximum at MR=0

Marginal revenue is a change in total revenue

TR=P*Q

=40Q-2Q^2

A change in function is found by differentiation

MR=dTR/dQ=40-4Q=0

4Q=40

Q=10

P=40-2*10

P=$20

the  price maximizes total expenditures for consumers is $20